# Scattering wave packets of hadrons in gauge theories: Preparation on a quantum computer

### v2

@@ III.1.1 Ground-state preparation
In practice, it is found that evolving the hopping term and the electric-field term for only one step, i.e., $N_{\rm GS}=1$, is enough to create |$\ket{\Omega}$ with high precision for the lattice size used in this work. One can thus identify $\left(\theta^{h},\theta^{\epsilon}\right)$ as the circuit parameters and perform optimization as instructed. The detailed circuit is depicted in Fig. 7. The quality of |$\ket{\Omega}$ arising from this VQE ansatz and optimization will be discussed in Sec. III.2.1. The cost of ground-state preparation only depends on the system size as $H_{n,n+1}^{h}$ is a nearest-neighbor term, effectively a 1-length meson. If the VQE demands $N_{\rm GS}>1$ layers, the CNOT-gate cost gets multiplied by $N_{\rm GS}$.
@@
III.1.2 Unitary implementation of creation operators
To implement $\mathcal{Q}_{\text{WP}}$, one can embed the non-unitary creation operators within a unitary operator that acts on an extended Hilbert space. In this work, the ancilla encoding introduced in Ref. Jordan et al. (2014) is used to realize such operators. Consider a non-unitary creation operator $b_{\Psi}^{\dagger}$. One can introduce an ancilla qubit, and define the following Hermitian operator:
@@ III.1.2 Unitary implementation of creation operators
As mentioned in Sec. II.2, the creation operator $b_{\Psi}^{\dagger}$ is a linear combination of many position-space mesonic creation operators with coefficients $C_{m,n}$, see Eq. (12) and subsequent equations. In general, these mesonic operators do not commute, and to circuitize the full creation operator, a Trotterization scheme is required. Nonetheless, the method above still works, up to a trotter error Childs et al. (2021), as long as the resulting |$\ket{\Psi}$ is properly normalized.
III.1.3 Singular-value-decomposition circuit
@@
Examples of circuit elements are explicitly shown in Fig. 8. The number of entangling gates in the $\mathcal{Q}_{\text{WP}}$ circuit depends on the meson length $|m-n|$, which determines the number of CNOT gates in $\mathcal{U}$ and $\mathcal{U}^{\dagger}$, as well as the size of the Jordan-Wigner Pauli string in $\mathcal{D}$. As discussed above, the wave-packet creation operator requires Trotterization due to non-commuting summands of mesonic operators. The Trotterization order, $N_{\text{Order}}$, and the number of Trotter steps, $N_{\text{Trotter}}$, are, therefore, introduced as circuit parameters. We consider a second-order product formula ($N_{\text{Order}}=2$) throughout this work, and the number of Trotter steps, $N_{\text{Trotter}}$, contributes multiplicatively toward the total resource counting. Furthermore, when the circuit is implemented on the hardware, we consider a truncation, $\theta_{c}$, on the coefficients $|C_{m,n}|$ when implementing $e^{-i\frac{\pi}{2}\Theta_{m,n}}$. This turns out to be effectively omitting the creation operators that produce larger mesons. As will be seen in Sec. III.2.2 and Appendix B, $\theta_{c}$ is chosen such that mostly 1-length mesons are created.
@@
The resource counting (number of qubits and entangling gates) of the circuits for ground-state, the full wave packet, and the 1-meson truncated wave-packet preparation are listed in Table 1. To further simplify the circuit, we have assumed that $N=4\mathbb{N}-2$ for positive integer $\mathbb{N}$ such that the Jordan-Wigner $\sigma^{\textbf{z}}$ strings in $\Theta_{m,n}$ are always implemented along the shorter path between $m$ and $n$. This choice is consistent with that from the original Jordan-Wigner transformation Eq. (II.1), as long as $(-1)^{Q+1}=1$ (see discussions regarding PBCs and Jordan-Wigner transform in Sec. II.1). Since we have restricted the analysis to the $Q=N/2$ sector, for $N=4\mathbb{N}-2$, such an efficient implementation is justified. Then for an $N$-site lattice, there are $N^{2}$ different $(m,n)$ pairs in position space. Recall that the ansatz only creates shorter mesons, and keep the other wrapping only when the forward-wrapped and backward-wrapped mesons have the same length. Thus, one can collect $N$ 0-length mesons and $2N$ $l$-length mesons where $0<l\leq\frac{N}{2}$. The creation operators of 0-length mesons are already diagonal, therefore require no $\mathcal{U}$ and $\mathcal{U}^{\dagger}$ circuits, and 2 CNOT gate for $e^{-i\theta\mathcal{D}}$. For each $l$-length meson creation operator, the number of entangling gates depends on $l$. In Fig. 8, one can see that $l+2$ and $42l+$ 4 CNOT gates are needed for $\mathcal{U}$ and $e^{-i\theta\mathcal{D}}$, respectively. Summing $l$ from 1 to $\frac{N}{2}$ and multiplying by the $2N$ factor, one can recover the $N^{3}$ dependence in Table 1. If the circuit is truncated such that only the $l\leq 1$ summands are considered, the gate count for individual operators will be constant, and the scaling of the total CNOT gates will be linear in $N$ as shown in Table 1.
Qubit count CNOT-gate count
$\mathcal{U}$ $\mathcal{D}$ Total
@@
Wave-packet circuit $12N+$ 1 $\frac{1}{4}N^{3}+\frac{5}{2}N^{2}$ $\frac{1}{2}N^{3}+5N^{2}+2N$ $\left(N^{3}+10N^{2}+2N\right)\times 2N_{\textrm{Trotter}}$
@@
III.2 Circuit implementation and results
In order to demonstrate how the algorithm of this work performs in practice, we present in this section hardware-implementation results for the wave-packet creation on a 6-site system in the $Z_{2}$ LGT in 1+1 D. These results are compared against the exact states obtained by classical simulations of the wave-packet circuit, referred to as ‘statevector’ evolution in the following. The $Z_{2}$ LGT parameters are chosen to be $m_{f}=1$ an $\epsilon=-0.3$, for which $b_{k}^{\dagger}$ creates momentum eigenstates with high fidelity (see Sec. II.3).
@@ III.2 Circuit implementation and results
III.2.1 Ground-state preparation
Figure 9: Probabilities of the physical basis states in the ground-state wave function. For 6 sites, there are 40 physical configurations in total. The $0^{\text{th}}$ state with the highest probability is the strong-coupling vacuum. The basis states are listed in Table 2 in Appendix E.
@@ III.2.1 Ground-state preparation
III.2.2 Full wave-packet simulation
@@
Figure 10: Physical basis-state probabilities of two wave packets in the case of $Z_{2}$ LGT with $m_{f}=1$ and $\epsilon=0.3$, generated on the Quantinuum H1-1 quantum computer compared with those obtained from |$\ket{\Psi}_{\text{exact}}$, |$\ket{\Psi}_{\text{ideal}}$, and |$\ket{\Psi}_{\text{trunc}}$, as defined in the text, as well as the associated local particle densities, $\chi_{n}$. The physical basis states are listed in Table 2 in Appendix E. The hardware results shown are after a symmetry-based error mitigation as discussed in the text. These agree reasonably well with the truncated-circuit output. The density plots clearly show the change in the shape of wave packets with varying $\sigma$, i.e., larger $\sigma$ results in a narrower wave packet in position space.
First note that there is a small but visible difference between the exact results and the ideal-circuit outputs. This discrepancy is attributed to a systematic error that originates from the unitary encoding in Sec. III.1.2. To ensure that the state |$\ket{\Psi}$ gets properly encoded, the conditions $b_{\Psi}\ket{\Omega}=0$ and $\left[b_{\Psi},b_{\Psi}^{\dagger}\right]=\mathds{1}$ need to hold. A small violation of these conditions, arising from small difference between optimized $b_{k}^{\dagger}$ and the exact ones, results in imperfect encoding of |$\ket{\Psi}$ using unitary operations. For an ideal creation operator ${b_{k}^{\dagger}}$ that satisfies the conditions stated, the circuit will prepare the wave-packet state assembled with $\Psi(k)\ket{k}$.
@@
Finally, the hardware results is contaminated by various sources of error prevalent to trapped-ion quantum devices Qua . Our resources in terms of both the number of shots and circuit depth forbid us from implementing proper error-mitigation protocols, such as the Pauli-twirling-based methods used in Ref. Farrell et al. (2023b). Still, it is possible to improve the outcome purely by post-processing the machine readout. An obvious consequence of the noisy hardware is a non-vanishing probability of leakage out of the physical Hilbert space, as the associated errors can violate gauge invariance. One can, therefore, consider a “symmetry-based” error mitigation scheme, which is to only count the events within the physical Hilbert space, and properly renormalize the final wave-function outcome. For wave packets in panels (a) and (b) of Fig. 10, respectively, 356 and 412 events out of the 500 shots are left within the physical Hilbert space after such an error-mitigation protocol, and 306 and 349 events have their ancilla measured to be 1.
@@ IV Summary and outlook
    We focused on simple local observables, such as basis-states probabilities, energy density, and local staggered charge density, to assess the quality of the wave packet generated on the hardware. In principle, other two- and $n$-point functions, including the non-local ones, need to be measured to fully verify the fidelity of the state produced. These, in turn, require additional quantum-circuit runs involving measurement of the desired operators, which were too costly given our limited computational credits on Quantinuum. Alternatively, one may resort to efficient entanglement-Hamiltonian tomography schemes Kokail et al. (2021a, b); Joshi et al. (2023); Mueller et al. (2023); Bringewatt et al. (2023), augmented by randomized measurement tools Ohliger et al. (2013); Elben et al. (2019); Huang et al. (2020); Vermersch et al. (2023); Elben et al. (2023), to reconstruct the system’s density matrix. These require many random measurements, and deemed expensive in our current study using Quantinuum’s resources. Such comprehensive measurement schemes, nonetheless, will be essential for the full scattering process, as deducing the nature of the post-collision state is an ultimate goal of the simulation.
@@
    State-of-the-art quantum hardware may soon reach capacity and capability needed to host larger systems. Therefore, wave-packet preparation in a confined gauge theory with a continuum limit, such as the $U(1)$, $SU(2)$, and $SU(3)$ LGTs in 1+1 D, should be feasible [see authors’ note below]. There is no qualitative difference in terms of circuit design, except bosonic operations are performed on a multi-qubit subspace for each gauge field. Various subcircuits for relevant operations, such as boson addition and subtraction, have, nonetheless, been developed in literature Shaw et al. (2020); Davoudi et al. (2022b) and can be straightforwardly ported to the algorithm of this work. Importantly, generating mesonic excitations in higher dimensional theories can likely proceed via the ansatz of this work, as mesons should still be well approximated by quark-antiquark pairs connected by an electric flux. Nonetheless, there will be more bare mesonic operators to account for different shapes and orientations of the fluxes in space.
@@
    While we have carefully assessed the algorithmic and statistical errors associated with our construct in the numerical benchmarks performed on small systems, it would be favorable to conduct a comprehensive analytical study of systematic errors introduced by the inexact exponentiation and Trotterization, see Sec. III.1. Such studies are common in the context of implementing the time-evolution operator within given accuracy (see e.g., Refs. Shaw et al. (2020); Kan and Nam (2021); Davoudi et al. (2022b) for gauge-theory examples) and need to be extended to state-preparation studies in gauge theories. Only such studies will enable accurate estimates of the resources required as a function of the desired wave-packet quality, and will allow comparing the efficiency of different methods in literature in near- and far-term eras of quantum computing. We defer such an analysis to future work.
@@
    While we did not discuss analog-simulation schemes in this work, we should remark that several interesting proposals have been developed in recent years to demonstrate analog simulators as a useful tool for studying scattering processes in simple gauge theories. For example, Ref. Surace and Lerose (2021) presents protocols to observe and measure selected meson-meson scattering processes in a $Z_{2}$ model in 1+1 D in Rydberg-atom arrays. Reference Belyansky et al. (2023) proposes a scheme to simulate collision of high-energy quark-antiquark and meson-meson wave packets in non-confining and confining regimes of the lattice Schwinger model with a topological $\theta$-term using circuit-QED platforms, after mapping the model to a bosonic equivalent. Furthermore, Ref. Su et al. (2024) proposes a particle-collision experiment in optical-superlattice quantum simulators of a lattice Schwinger model with a $\theta$-term. It would be interesting to see how the mesonic ansatz of this work can facilitate these and future analog and hybrid analog-digital simulations of scattering processes in various LGTs.
@@
    The ultimate goal is to develop expressive and efficient forms of hadronic wave-packet ansatze, particularly for baryons and nuclei, relevant to quantum chromodynamics. Insights from classical LGT studies of hadrons and nuclei will likely be crucial in coming up with such ansatze. In particular, a wealth of results on the spectrum Detmold et al. (2019); Bulava et al. (2022); Prelovsek (2023) and structure Constantinou et al. (2021); Hagler (2010) of hadrons and nuclei using lattice-QCD computations may need to be incorporated in state-preparation schemes on quantum computers to reduce the required quantum resources. These results can potentially both inform the form of the hadronic ansatz, and aid in parameter optimizations. The strategy of this work can be generalized to take advantage of such opportunities.
Authors’ note. In the final stage of drafting this manuscript, another manuscript on wave-packet preparation in the Schwinger model was released Farrell et al. (2024). This reference prepares the wave packet in the interacting theory using an improved VQE algorithm that finds a low-depth quantum circuit which creates the wave packets out of the interacting vacuum. The interacting vacuum itself is prepared using the improved VQE developed by the authors. The optimization for the VQE is performed by comparing against a wave packet that is adiabatically produced out of a bare mesonic wave packet with the aid of classical computing. Using such a hybrid classical-quantum strategy, and performing state-of-the-art noise mitigation techniques suited to the IBM quantum-computing platform, the authors create and evolve single wave packets on lattices as large as 112 staggered sites (112 qubits). We leave comparing the quality of wave packets prepared via this strategy and that presented in our work to future studies.

### v3

@@ 3.1.1 Ground-state preparation
In practice, it is found that evolving the hopping term and the electric-field term for only one step, i.e., $N_{\rm GS}=1$, is enough to create |$\ket{\Omega}$ with high precision for the lattice size used in this work. One can thus identify $\left(\theta^{h},\theta^{\epsilon}\right)$ as the circuit parameters and perform optimization as instructed. The detailed circuit is depicted in Fig. 7. The quality of |$\ket{\Omega}$ arising from this VQE ansatz and optimization will be discussed in Sec. 3.2.1. The cost of ground-state preparation only depends on the system size as $H_{n,n+1}^{h}$ is a nearest-neighbor term, effectively a 1-length meson. If the VQE demands $N_{\rm GS}>1$ layers, the CNOT-gate cost gets multiplied by $N_{\rm GS}$.
@@
3.1.2 Unitary implementation of creation operators
To implement $\mathcal{Q}_{\text{WP}}$, one can embed the non-unitary creation operators within a unitary operator that acts on an extended Hilbert space. In this work, the ancilla encoding introduced in Ref. [99] is used to realize such operators. Consider a non-unitary creation operator $b_{\Psi}^{\dagger}$. One can introduce an ancilla qubit, and define the following Hermitian operator:
@@ 3.1.2 Unitary implementation of creation operators
As mentioned in Sec. 2.2, the creation operator $b_{\Psi}^{\dagger}$ is a linear combination of many position-space mesonic creation operators with coefficients $C_{m,n}$, see Eq. (12) and subsequent equations. In general, these mesonic operators do not commute, and to circuitize the full creation operator, a Trotterization scheme is required. Nonetheless, the method above still works, up to a trotter error [128], as long as the resulting |$\ket{\Psi}$ is properly normalized.
3.1.3 Singular-value-decomposition circuit
@@
Examples of circuit elements are explicitly shown in Fig. 8. The number of entangling gates in the $\mathcal{Q}_{\text{WP}}$ circuit depends on the meson length $|m-n|$, which determines the number of CNOT gates in $\mathcal{U}$ and $\mathcal{U}^{\dagger}$, as well as the size of the Jordan-Wigner Pauli string in $\mathcal{D}$. As discussed above, the wave-packet creation operator requires Trotterization due to non-commuting summands of mesonic operators. The Trotterization order, $N_{\text{Order}}$, and the number of Trotter steps, $N_{\text{Trotter}}$, are, therefore, introduced as circuit parameters. We consider a second-order product formula ($N_{\text{Order}}=2$) throughout this work, and the number of Trotter steps, $N_{\text{Trotter}}$, contributes multiplicatively toward the total resource counting. Furthermore, when the circuit is implemented on the hardware, we consider a truncation, $\theta_{c}$, on the coefficients $|C_{m,n}|$ when implementing $e^{-i\frac{\pi}{2}\Theta_{m,n}}$. This turns out to be effectively omitting the creation operators that produce larger mesons. To what extent these large mesons are truncated depends on the parameters of the theory ($m_{f}$, $\epsilon$) and of the wave packet ($\sigma$, $\mu$ and $k_{0}$), which will be further discussed in Appendix B. A general observation is that the distribution of $|C_{m,n}|$ tends to be more uniform across sites $(m,n)$ in the low-mass regime. This implies more resources are required to prepare the state due to a higher surviving chance of the mesonic operators at a given $\theta_{c}$. As will be seen in Sec. 3.2.2 and Appendix B, $\theta_{c}$ is chosen such that mostly 1-length mesons are created.
@@
The resource counting (number of qubits and entangling gates) of the circuits for ground-state, the full wave packet, and the 1-meson truncated wave-packet preparation are listed in Table 1. To further simplify the circuit, we have assumed that $N=4\mathbb{N}-2$ for positive integer $\mathbb{N}$ such that the Jordan-Wigner $\sigma^{\textbf{z}}$ strings in $\Theta_{m,n}$ are always implemented along the shorter path between $m$ and $n$. This choice is consistent with that from the original Jordan-Wigner transformation Eq. (3), as long as $(-1)^{Q+1}=1$ (see discussions regarding PBCs and Jordan-Wigner transform in Sec. 2.1). Since we have restricted the analysis to the $Q=N/2$ sector, for $N=4\mathbb{N}-2$, such an efficient implementation is justified. Then for an $N$-site lattice, there are $N^{2}$ different $(m,n)$ pairs in position space. Recall that the ansatz only creates shorter mesons, and keep the other wrapping only when the forward-wrapped and backward-wrapped mesons have the same length. Thus, one can collect $N$ 0-length mesons and $2N$ $l$-length mesons where $0<l\leq\frac{N}{2}$. The creation operators of 0-length mesons are already diagonal, therefore require no $\mathcal{U}$ and $\mathcal{U}^{\dagger}$ circuits, and 2 CNOT gate for $e^{-i\theta\mathcal{D}}$. For each $l$-length meson creation operator, the number of entangling gates depends on $l$. In Fig. 8, one can see that $l+2$ and $42l+$ 4 CNOT gates are needed for $\mathcal{U}$ and $e^{-i\theta\mathcal{D}}$, respectively. Summing $l$ from 1 to $\frac{N}{2}$ and multiplying by the $2N$ factor, one can recover the $N^{3}$ dependence in Table 1. If the circuit is truncated such that only the $l\leq 1$ summands are considered, the gate count for individual operators will be constant, and the scaling of the total CNOT gates will be linear in $N$ as shown in Table 1. In summary, the number of gates required to prepare the wave packet scales at worst as $N^{3}\times N_{\rm Trotter}$, and can be reduced to $N\times N_{\rm Trotter}$ if only mesons with length less or equal to 1 are implemented.
Overall, the performance of the algorithm is also a function of the performance of the ansatz, which as observed in Fig. 5, depends on the systems parameters $N$, $m_{f}$, and $\epsilon$. Unfortunately, without the analytical knowledge of the states (as is the case in nonperturbative theories), an analytic bound cannot be found on the performance of the ansatz, hence the algorithm. Our study, nonetheless, establishes empirical evidence of an acceptable algorithm performance in small systems and for a range of parameters.
Qubit CNOT-gate count
count $\mathcal{U}$ $\mathcal{D}$ Total
@@
Wave-packet circuit $12N+$ 1 $\frac{1}{4}N^{3}+\frac{5}{2}N^{2}$ $\frac{1}{2}N^{3}+5N^{2}+2N$ $\left(N^{3}+10N^{2}+2N\right)$
(full) ×$\times 2N_{\textrm{Trotter}}$
@@
3.2 Circuit implementation and results
In order to demonstrate how the algorithm of this work performs in practice, we present in this section hardware-implementation results for the wave-packet creation on a 6-site system in the $Z_{2}$ LGT in 1+1 D. These results are compared against the exact states obtained by classical simulations of the wave-packet circuit, referred to as ‘statevector’ evolution in the following. The $Z_{2}$ LGT parameters are chosen to be $m_{f}=1$ an $\epsilon=-0.3$, for which $b_{k}^{\dagger}$ creates momentum eigenstates with high fidelity (see Sec. 2.3).
@@ 3.2 Circuit implementation and results
3.2.1 Ground-state preparation
Figure 9: Probabilities of the physical basis states in the ground-state wave function. For 6 sites, there are 40 physical configurations in total. (Classically-simulated) VQE is able to approach the exact-diagonalization results with high accuracy. The $0^{\text{th}}$ state with the highest probability is the strong-coupling vacuum. The basis states are listed in Table 2 in Appendix E.
@@ 3.2.1 Ground-state preparation
3.2.2 Full wave-packet simulation
@@
Figure 10: Physical basis-state probabilities of two wave packets in the case of $Z_{2}$ LGT with $m_{f}=1$ and $\epsilon=-0.3$, generated on the Quantinuum H1-1 quantum computer compared with those obtained from |$\ket{\Psi}_{\text{exact}}$, |$\ket{\Psi}_{\text{ideal}}$, and |$\ket{\Psi}_{\text{trunc}}$, which are wave-packet states from exact optimization, and from classical simulation of exact and truncated circuits, respectively, as defined in the text. The associated local particle densities, $\chi_{n}$ calculated based on these states, along with that obtained from hardware, are presented on the right panels. The physical basis states are listed in Table 2 in Appendix E. The hardware results shown are after a symmetry-based error mitigation as discussed in the text. These agree reasonably well with the truncated-circuit output. The density plots clearly show the change in the shape of wave packets with varying $\sigma$, i.e., larger $\sigma$ results in a narrower wave packet in position space.
First note that there is a small but visible difference between the exact results and the ideal-circuit outputs. This discrepancy is attributed to a systematic error that originates from the unitary encoding in Sec. 3.1.2. To ensure that the state |$\ket{\Psi}$ gets properly encoded, the conditions $b_{\Psi}\ket{\Omega}=0$ and $\left[b_{\Psi},b_{\Psi}^{\dagger}\right]=\mathds{1}$ need to hold. A small violation of these conditions, arising from small difference between optimized $b_{k}^{\dagger}$ and the exact ones, results in imperfect encoding of |$\ket{\Psi}$ using unitary operations. For an ideal creation operator ${b_{k}^{\dagger}}$ that satisfies the conditions stated, the circuit will prepare the wave-packet state assembled with $\Psi(k)\ket{k}$.
@@
Finally, the hardware results is contaminated by various sources of error prevalent to trapped-ion quantum devices [129]. Our resources in terms of both the number of shots and circuit depth forbid us from implementing proper error-mitigation protocols, such as the Pauli-twirling-based methods used in Ref. [125]. Still, it is possible to improve the outcome purely by post-processing the machine readout. An obvious consequence of the noisy hardware is a non-vanishing probability of leakage out of the physical Hilbert space, as the associated errors can violate gauge invariance. One can, therefore, consider a “symmetry-based” error mitigation scheme, which is to only count the events within the physical Hilbert space, and properly renormalize the final wave-function outcome. For wave packets in panels (a) and (b) of Fig. 10, respectively, 356 and 412 events out of the 500 shots are left within the physical Hilbert space after such an error-mitigation protocol, and 306 and 349 events have their ancilla measured to be 1.
@@ 4 Summary and outlook
The algorithm of this work, therefore, provides an efficient and effective procedure to create hadronic wave-packet states for scattering in (1+1)D gauge theories. While ground and low-energy state preparation in such models is feasible with classical methods, such as tensor networks, of particular interest are scattering processes at high energies with abundant entanglement generation, which are challenging for classical algorithms. Hence, combining an efficient state-preparation algorithm, such as that provided in this work, and time-evolution and measurement steps on a quantum computer, can facilitate studies of scattering on quantum hardware in the future.
@@
    We focused on simple local observables, such as basis-states probabilities, energy density, and local staggered charge density, to assess the quality of the wave packet generated on the hardware. In principle, other two- and $n$-point functions, including the non-local ones, need to be measured to fully verify the fidelity of the state produced. These, in turn, require additional quantum-circuit runs involving measurement of the desired operators, which were too costly given our limited computational credits on Quantinuum. Alternatively, one may resort to efficient entanglement-Hamiltonian tomography schemes [130, 131, 132, 46, 133], augmented by randomized measurement tools [134, 135, 136, 137, 138], to reconstruct the system’s density matrix. These require many random measurements, and deemed expensive in our current study using Quantinuum’s resources. Such comprehensive measurement schemes, nonetheless, will be essential for the full scattering process, as deducing the nature of the post-collision state is an ultimate goal of the simulation.
@@
    State-of-the-art quantum hardware may soon reach capacity and capability needed to host larger systems. Therefore, wave-packet preparation in a confined gauge theory with a continuum limit, such as the $U(1)$, $SU(2)$, and $SU(3)$ LGTs in 1+1 D, should be feasible [see authors’ note below]. There is no qualitative difference in terms of circuit design, except bosonic operations are performed on a multi-qubit subspace for each gauge field. Various subcircuits for relevant operations, such as boson addition and subtraction, have, nonetheless, been developed in literature [50, 54] and can be straightforwardly ported to the algorithm of this work. Importantly, generating mesonic excitations in higher dimensional theories can likely proceed via the ansatz of this work, as mesons should still be well approximated by quark-antiquark pairs connected by an electric flux. Nonetheless, there will be more bare mesonic operators to account for different shapes and orientations of the fluxes in space.
@@
    While we have carefully assessed the algorithmic and statistical errors associated with our construct in the numerical benchmarks performed on small systems, it would be favorable to conduct a comprehensive analytical study of systematic errors introduced by the inexact exponentiation and Trotterization, see Sec. 3.1. Such studies are common in the context of implementing the time-evolution operator within given accuracy (see e.g., Refs. [50, 53, 54] for gauge-theory examples) and need to be extended to state-preparation studies in gauge theories. Only such studies will enable accurate estimates of the resources required as a function of the desired wave-packet quality, and will allow comparing the efficiency of different methods in literature in near- and far-term eras of quantum computing. We defer such an analysis to future work.
@@
    While we did not discuss analog-simulation schemes in this work, we should remark that several interesting proposals have been developed in recent years to demonstrate analog simulators as a useful tool for studying scattering processes in simple gauge theories. For example, Ref. [87] presents protocols to observe and measure selected meson-meson scattering processes in a $Z_{2}$ model in 1+1 D in Rydberg-atom arrays. Reference [34] proposes a scheme to simulate collision of high-energy quark-antiquark and meson-meson wave packets in non-confining and confining regimes of the lattice Schwinger model with a topological $\theta$-term using circuit-QED platforms, after mapping the model to a bosonic equivalent. Furthermore, Ref. [139] proposes a particle-collision experiment in optical-superlattice quantum simulators of a lattice Schwinger model with a $\theta$-term. It would be interesting to see how the mesonic ansatz of this work can facilitate these and future analog and hybrid analog-digital simulations of scattering processes in various LGTs.
@@
    The ultimate goal is to develop expressive and efficient forms of hadronic wave-packet ansatze, particularly for baryons and nuclei, relevant to quantum chromodynamics. Insights from classical LGT studies of hadrons and nuclei will likely be crucial in coming up with such ansatze. In particular, a wealth of results on the spectrum [140, 141, 142] and structure [143, 144] of hadrons and nuclei using lattice-QCD computations may need to be incorporated in state-preparation schemes on quantum computers to reduce the required quantum resources. These results can potentially both inform the form of the hadronic ansatz, and aid in parameter optimizations. The strategy of this work can be generalized to take advantage of such opportunities.
Authors’ note. In the final stage of drafting this manuscript, another manuscript on wave-packet preparation in the Schwinger model was released [145]. This reference prepares the wave packet in the interacting theory using an improved VQE algorithm that finds a low-depth quantum circuit which creates the wave packets out of the interacting vacuum. The interacting vacuum itself is prepared using the improved VQE developed by the authors. The optimization for the VQE is performed by comparing against a wave packet that is adiabatically produced out of a bare mesonic wave packet with the aid of classical computing. Using such a hybrid classical-quantum strategy, and performing state-of-the-art noise mitigation techniques suited to the IBM quantum-computing platform, the authors create and evolve single wave packets on lattices as large as 112 staggered sites (112 qubits). We leave comparing the quality of wave packets prepared via this strategy and that presented in our work to future studies.